OpenStudy (anonymous):

Given that sin equals 2/7 what is the value of tan?

OpenStudy (anonymous):



OpenStudy (anonymous):

there is the picture on an angle \(\theta\) where \(\sin(\theta)=\frac{2}{7}\) what you need is the third side of the triangle, which you get by pythagoras

OpenStudy (anonymous):

There was no drawing and in a test. I realy need the answer

OpenStudy (anonymous):

i made the drawing so you could see what to do to solve this what you need is the third side of the triangle \[a^2+b^2=h^2\] \[a^2+2^2=7^2\] \[a^2=7^2-2^2=49-4=45\] and so \(a=\sqrt{45}=3\sqrt{5}\)

OpenStudy (anonymous):

But how do i get tan?